Calculate the de Broglie wavelength associated with a moving bullet, whose mass is 100 gm and velocity is 360 km per hour.
1
Expert's answer
2020-11-19T09:29:58-0500
"\lambda=\frac{h}{mv}\\\lambda=\frac{6.626\cdot10^{-34}}{0.1\frac{360}{3.6}}=6.626\cdot10^{-35}\ m"